OpenStudy (anonymous):

Write the equation of the line which passes through (-4, 3) and with slope of 1 in slope-intercept form. y=x+3 3x−4y=1 y−3=x+4 y=x+7

OpenStudy (anonymous):

slope intercept form : y = mx + b slope(m) = 1 (-4,3) x = -4 and y = 3 now we sub and find b 3 = 1(-4) + b 3 = -4 + b 3 + 4 = b 7 = b so your equation is : y = x + 7

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

i have one last problem like this i need help with, could you help me with it ?

OpenStudy (anonymous):

sure...what ya got ?

OpenStudy (anonymous):

Write the equation of the line which passes through (5, -2) and (3, 4) in standard form.

OpenStudy (anonymous):

ok...first we need to find the slope by using the slope formula. slope(m) = (y2 - y1) / (x2 - x1) (5,-2) x1 = 5 and y1 = -2 (3,4) x2 = 3 and y2 = 4 now we sub slope(m) = (4 - (-2) / (3 - 5) slope(m) = (4 + 2) / -2 slope(m) = 6/-2 which = - 3 now we will use y = mx + b and solve for b slope(m) = -3 you can use either of your set of points (3,4) x = 3 and y = 4 now we sub 4 = -3(3) + b 4 = -9 + b 4 + 9 = b 13 = b equation is : y = -3x + 13 However, we need it in standard form Ax + By = C y = -3x + 13 --- add 3x to both sides 3x + y = 13 ====> standard form equation

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

nope! thanks so much!!!

OpenStudy (anonymous):

anytime :)